a trail mix that costs $2.45 per pound is mixed with a trail mix that costs $2.30 per pound. how much of each type of trail mix must be used to have 30 pounds of a trail mix that costs $2.35 per pound
Thanks

Answer by checkley79(3341)

2.45x+2.30(30-x)=2.35*30
2.45x+69-2.30x=70.5
.15x=70.5-69
.15x=1.5
x=1.5/.15
x=10 pounds of the $2.45 mix is used.
30-10=20 pounds of the 42.30 mix ix used.
Proof:
2.45*10+2.30*20=2.35*30
24.5+46=70.5
70.5=70.5